# FirstReferences

### v2

@@ Abstract
Recently, the high degree of control and tunability of programmable quantum platforms—such as trapped ions [8], superconducting qubits [12], and photonic devices [13]—potentially provides an alternative approach to large-scale classical computations for studying condensed-phase chemical dynamics through direct quantum simulation. In particular, trapped-ion analog and analog-digital simulators have already realized a variety of chemical dynamics with fully programmable system parameters and accurate time-resolved features [14, 15, 16, 17, 18]. With engineered reservoirs on their native spin and bosonic degrees of freedom, trapped-ion systems have also been recently used to study non-equilibrium quantum reaction dynamics [19, 20, 21, 22].
@@
In this work, we present the first trapped-ion simulation in which unitary spin–phonon couplings and mode-selective dissipation via cooling are independently programmed, enabling real-time open-system emulation of the excitation transfer dynamics in a two-mode LVCM. Using both ground-state and optical qubits, we achieve simultaneous control over mode frequencies, vibronic couplings, and dissipation rates. We characterize the low-temperature transfer rate across a wide parameter range, highlighting the roles of mode degeneracy and vibronic coupling strength. We focus on two transfer regimes (see Fig. 1A): charge transfer (CT) and vibrationally assisted exciton transfer (VAET), which correspond to strong and weak vibronic coupling, respectively  [^1] We note that charge transfer can also occur in the weak vibronic coupling regime. However, in this study, we use this process to refer specifically to the strong vibronic coupling regime, with which it is more commonly associated..
Figure 1: Quantum simulation of a two-mode LVCM with engineered reservoirs. (A) Schematic diagram illustrating two regimes of transfer dynamics in the LVCM, defined by vibronic coupling strengths. In the VAET regime, the harmonic oscillators are weakly coupled to the donor–acceptor electronic system, whereas the strong vibronic couplings in the CT regime introduce significant displacements, resulting in distinct two-dimensional potential energy surfaces for the donor (red) and acceptor (blue) states. The electronic coupling opens a $2V$ avoided crossing between these surfaces. For clarity, we show only the $y_{2}=0$ cut across the two-dimensional potential energy surfaces, with the vibrational levels associated with the fast $(\omega_{1})$ and slow ($\omega_{2}$) modes, represented by their wavefunctions and horizontal lines, respectively. In both regimes, the vibrational modes undergo dissipation, represented by the wiggly arrows. (B) Experimental setup for studying open-system LVCM dynamics with two vibrational modes using a 171Yb+ -172Yb+ ion chain, whose motional degrees of freedom are shared (represented by the connecting spring). Insets: simplified level schemes for 171Yb+ and 172Yb+ qubits. Stimulated Raman transitions on the 171Yb+ ground-state qubit with 355 nm beams (purple) are used to engineer the Hamiltonian in Eq. (1). The optical qubit of 172Yb+ is addressed with a 435.5 nm beam (light blue) and a 935 nm repumper beam (red line in the 172Yb+ inset) for sympathetic cooling. (C) Experimental sequence used to measure the time-resolved evolution of the excitation at the donor electronic site (see Methods for details). Each dashed red line indicates the state of the system after each preparation pulse.
@@
where $\sigma_{x,z}$ are the Pauli operators acting on the donor and acceptor electronic sites, |$\ket{D}\equiv\ket{\uparrow}_{z}$ and |$\ket{A}\equiv\ket{\downarrow}_{z}$, with an energy difference $\Delta E$ and an electronic coupling strength $V$. For $\Delta E>0$, the Hamiltonian describes an exothermic reaction when the excitation is transferred from the donor site to the acceptor site. Each harmonic oscillator $i$ with a vibrational energy $\omega_{i}$ is associated with creation ($a_{i}^{\dagger}$) and annihilation ($a_{i}$) operators, and it is linearly coupled to the electronic sites at a rate $g_{i}$.
@@
In this model, the charge transfer (CT) regime is realized when the vibronic coupling is comparable to or larger than the characteristic harmonic frequency scale ($g_{i}\gtrsim\omega_{i}$, also known as strong coupling in quantum optics), which is an attribute of many situations involving charge motion in chemical and biological reactions ranging from redox catalysis to solvent-induced electronic delocalization [25, 26, 27, 28]. The donor and acceptor electronic sites in this regime are described by uncoupled two-dimensional potential energy surfaces with respect to the ($y_{1}$, $y_{2}$) spatial coordinates, where $y_{i}=y_{i0}(a_{i}+a_{i}^{\dagger})/2$ with $y_{i0}=\sqrt{1/2m\omega_{i}}$ and $m$ being the particle mass. The electronic coupling $V$ mixes the two potential energy surfaces by opening an avoided crossing with a gap of $2V$ at their intersection. A strong vibronic coupling ($g_{i}\gtrsim\omega_{i}$) distorts the energy landscape by inducing a displacement of $g_{i}/\omega_{i}$ between the donor and acceptor potential energy surfaces along $y_{i}$, providing the electronic coupling with a dependence on the overlaps of the displaced oscillator states [29, 28, 20]. During CT, the donor population undergoes crossing of the energy barrier along an effective reaction coordinate, roughly defined by the axis $y_{1}+y_{2}$ in the degenerate ($\omega_{1}=\omega_{2}$) case. The total reorganization energy of the system is defined as $\lambda=\sum_{i}\lambda_{i}\equiv\sum_{i}g_{i}^{2}/\omega_{i}$, which is the energy required to displace the wave packet on an uncoupled potential energy surface by the distances of $|g_{1}/\omega_{1}|$ and $|g_{2}/\omega_{2}|$ along the $y_{1}$ and $y_{2}$ axes, respectively. When the electronic coupling is small ($V\approx 0$), the reorganization energy $\lambda$ and energy difference between the two surfaces $\Delta E$ determine the classical activation energy, defined by $U=(\Delta E+\lambda)^{2}/4\lambda$ [28, 20].
@@
Conversely, a weak vibronic coupling ($g_{i}\ll\omega_{i}$) is more characteristic of vibrationally assisted exciton transfer (VAET), as occurs between the pigments in light-harvesting compounds and their reaction centers [14]. In this regime, the excitation transfer between the donor and acceptor sites with an energy separation $\Delta E$ is enabled by the electronic coupling term $V\sigma_{x}$, resulting in the two eigenenergies of the total electronic system differing by $\sqrt{\Delta E^{2}+(2V)^{2}}$. Different from the CT regime, where the strong vibronic couplings displace the donor and acceptor energy surfaces, thereby defining the vibronic states that participate in the excitation transfer, the couplings between the electronic system and the vibrational modes in the VAET regime are weak. These weak couplings facilitate quantized energy exchange between the oscillators and the excitation sites, leading to transfer resonances at $\Delta E\approx\sqrt{(\ell_{1}\omega_{1}+\ell_{2}\omega_{2})^{2}-(2V)^{2}}$, with $\ell_{1}$ and $\ell_{2}$ being integers. At resonance, the combined vibrational energy $\ell_{1}\omega_{1}+\ell_{2}\omega_{2}$ provided by the two modes exactly bridges the electronic energy gap, thereby assisting the transfer [14, 21, 30].
@@
In most realistic chemical situations, the interaction between the donor-acceptor system and the external environment causes the vibrational modes to undergo incoherent dissipation. Under the assumptions of linear coupling between the vibrational modes and a Markovian environment with Ohmic spectral densities [31, 28] and that the vibrational dissipation rate is weaker than the bath vibrational and thermal energies ($\gamma_{i}\ll\omega_{i},k_{B}T_{i}$), the dissipative dynamics of the system can be effectively described by a Lindblad master equation:
@@
Figure 2: Transfer rates in the CT regime. (A) Transfer rate spectrum of degenerate CT ($\omega_{1}\!=\!\omega_{2}\!\equiv\!\omega$) with $(V,g_{1},\gamma_{1},g_{2},\gamma_{2})=(0.200,1.200,0.036,1.100,0.040)\omega$. Red circles represent experimental data with error bars estimated via bootstrapping (see Methods). The solid red curve shows the transfer rate calculated from Eq. (2), using the definition in Eq. (3) and including spin decoherence ($\gamma_{z}\!=\!0.0014\omega$) and motional dephasing of both modes ($\gamma_{im}\!=\!0.0040\omega$, with $i=1,2$). The blue curve shows the numerical result for the single-mode CT case, where $\omega_{2}\!=\!g_{2}\!=\!\gamma_{2}\!=\!\gamma_{2m}\!=\!0$. (B and C) Experimental and numerical donor population evolution $P_{D}(t)$ versus energy gap $\Delta E$ and the number of vibrational oscillations $\omega t/2\pi$, using the same parameters as in the red circles and solid red curve in (A), respectively. Here, the detunings from the two tilt modes for encoding the degenerate vibrational energies are both set to $\delta_{1}=\delta_{2}=-2\pi\times 5$ kHz. (D to F) Same layout as in (A to C) for the non-degenerate CT case ($\omega_{1}\!>\!\omega_{2}$) with $(V,g_{1},\gamma_{1},\omega_{2},\gamma_{z},\gamma_{im})=(0.138,1.029,0.023,0.375,0.0009,0.0025)\omega_{1}$ and $(g_{2},\gamma_{2})\!=\!(0.993,0.027)\omega_{2}$. The characteristic timescale is defined by the fast vibrational frequency $\omega_{1}$. In this case, the detuning from the inner (outer) tilt mode, which encodes the fast (slow) vibrational energy, is set to $\delta_{1}\!=\!-2\pi\!\times\!8$ kHz ($\delta_{2}\!=\!-2\pi\!\times\!3$ kHz).
To experimentally realize the multi-mode LVCM, we use a dual-species chain of one 171Yb+ ion and one 172Yb+ ion trapped in a linear Paul trap [20]. We encode the electronic degree of freedom in the two hyperfine clock states of the 171Yb+ ground-state qubit, |$\ket{{}^{2}S_{1/2},F=1,m_{F}=0}\equiv\ket{\uparrow}_{z}$ and |$\ket{{}^{2}S_{1/2},F=0,m_{F}=0}\equiv\ket{\downarrow}_{z}$ with the frequency splitting of $\omega_{\text{hf}}=2\pi\times 12.642$ GHz (see Fig. 1B). In this work, we use the two radial out-of-phase modes (also referred to as tilt modes), each along an orthogonal radial principal trap axis, to represent $\omega_{1}$ and $\omega_{2}$ (see Methods).
Figure 3: Transitioning from CT to VAET ($\omega_{1}\!>\!\omega_{2}$). (A) Transfer rate spectrum for $(V,g_{1},\gamma_{1},\omega_{2})=(0.063,1.288,0.015,0.375)\omega_{1}$ and $(g_{2},\gamma_{2})=(0.607,0.067)\omega_{2}$. Red circles represent experimental data with error bars estimated via bootstrapping (see Methods). The solid red curve shows the transfer rate calculated from Eq. (2), using the definition in Eq. (3) and including spin decoherence ($\gamma_{z}=0.0009\omega_{1}$) and motional dephasing of both modes ($\gamma_{im}=0.0025\omega_{1}$, with $i=1,2$). The blue curve shows the numerical result for the single-mode case, where $\omega_{2}=g_{2}=\gamma_{2}=\gamma_{2m}=0$. (B and C) Experimental and numerical donor population evolution $P_{D}(t)$ versus energy gap $\Delta E$ and the number of vibrational oscillations of the fast mode $\omega_{1}t/2\pi$, using the same parameters as in the red circles and solid red curve in (A), respectively. Here, the detuning from the inner (outer) tilt mode, which encodes the fast (slow) vibrational energy, is set to $\delta_{1}=-2\pi\times 8$ kHz ($\delta_{2}=-2\pi\times 3$ kHz).
Our experiment is performed in a driven rotating frame, where we use two $\pi/2$ pulses to map the $z$ spin basis in Eq. (1) onto the $y$ basis. This method allows us to use the ion-light interactions between the 171Yb+ qubit and the 355 nm Raman laser tones to independently engineer individual terms of the Hamiltonian in Eq. (1) with precise control. Two laser tones resonant with the qubit frequency generate the single-qubit operations ($V\sigma_{x}$ and $({\Delta E}/{2})\,\sigma_{z}$ terms), while two other laser tone pairs at frequencies ±$\pm\mu_{i}=\pm(\omega_{\text{tilt},i}+\delta_{i})$ from the qubit resonance realize the vibronic coupling and harmonic terms in Eq. (1). Here, $\delta_{i}$ is the detuning from its respective radial tilt mode at frequency $\omega_{\text{tilt},i}$, which determines the vibrational energy $\omega_{i}$ through $\delta_{i}\equiv-\omega_{i}$ [32, 20].
The engineered dissipation of the vibrational modes is realized by driving the narrow transitions from the ground-state manifold |$\ket{g}\!\equiv\!\ket{{}^{2}S_{1/2}}$ to the optical metastable-state manifold |$\ket{o}\equiv\ket{{}^{2}D_{3/2}}$ of the 172Yb+ ion, detuned by $-\omega_{\text{tilt},i}$ from resonance, using a total of four laser tones to address all the involved Zeeman sub-levels. Together with a 935 nm repumper beam to optically reset the electronic excitation to the ground states, this results in sympathetic cooling on both radial tilt modes of the chain with independently tunable dissipation rates. This is equivalent to generating a structured bath of continuous harmonic oscillators with two Lorentzian spectral densities centered at $\omega_{1}$ and $\omega_{2}$, with full widths at half maximum $\gamma_{1}$ and $\gamma_{2}$, respectively [33, 21] (see Appendix D). In this setup, all the system parameters—including the bath properties—are determined by the frequency and power of the laser tones used to generate the corresponding ion-light interactions. Therefore, they can be precisely tuned and independently calibrated [20], with the magnitudes of the system parameters used in our study being much larger than the decoherence rates caused by experimental imperfections (see Appendix B). The experimental sequence is discussed in Methods and summarized in Fig. 1C.
In this work, we focus on the non-perturbative quantum regime [28], where the electronic coupling strength is strong ($|V|\!\sim\!\lambda_{i}/4$), and the bath temperatures are low ($\bar{n}_{i}\sim 0.1$–0.2, see Methods). We compare the excitation transfer behaviors of the two-mode models with those of their single-mode counterparts. In the following, we characterize the transfer dynamics by measuring the inverse lifetime of the donor population ($P_{D}=(\langle\sigma_{z}\rangle+1)/2$), which is described by [29, 28, 20, 34]:
@@
Charge transfer (CT) - We first study non-perturbative CT with $g_{i}\sim\omega_{i}$ in two different cases: degenerate ($\omega_{1}=\omega_{2}\equiv\omega$) and non-degenerate ($\omega_{1}>\omega_{2}$). At large electronic coupling $|V|\sim\lambda_{i}/4$, the avoided crossing becomes appreciable, which leads to the hybridization of the donor and acceptor potential energy surfaces into upper and lower adiabatic energy surfaces, resulting in eigenstates that are superpositions of donor and acceptor vibronic states [28, 20]. In the degenerate case, we observe in Fig. 2A that, in the two-mode model, the exothermic region with monotonically increasing transfer rates—the cumulative broadening effect of strong electronic coupling and weak dissipation rates ($|V|>\gamma_{i}$)—extends to a higher energy gap, $\Delta E\approx 4\omega$, compared to its single-mode counterpart ($\Delta E\approx 3\omega$). This extension can be explained by the larger phase-space volume of the adiabatic energy surfaces when there are two modes available rather than just one.
@@
Moreover, we remark that resonant peaks at $\Delta E\approx\ell\omega$, with $\ell$ being an integer, appear when the energy difference is sufficiently high ($\Delta E\gtrsim 5\omega$). Under this energy gap condition, the initially localized donor state has significant overlaps with the eigenstates of the upper hybridized surface, leading to population trapping, as explained in Refs. [28, 20]. At these high resonances, the trapped population is released from the upper hybridized surface to the lower hybridized surface during the evolution. Due to the large energy offset and the increased phase-space volume in the two-mode case compared to the single-mode case, the steady state has greater overlaps with the acceptor states, as shown in Figs. 2B and 2C. This results in enhanced transfer rates at large energy gaps $\Delta E\gtrsim 5\omega$.
@@
Conversely, when $\omega_{1}>\omega_{2}$, the donor-acceptor energy landscape supports highly delocalized states along $y_{2}$, as the system is deeply in the adiabatic regime ($|V|>\lambda_{2}/4$) [28, 20]. This introduces multiple states delocalized along the $y_{2}$ direction, in addition to the existing delocalization along $y_{1}$ provided by the fast vibrational mode levels (see Fig. 1A), thereby enabling additional transfer channels among the highly delocalized two-dimensional vibronic states. Thus, the transfer process in the non-degenerate case is less sensitive to the donor-acceptor energy offset, making it more robust to variations in this parameter. Due to our experimental resolution (see Appendix B), we cannot resolve the resonances associated with these transfer channels and instead observe a smooth transfer rate curve in Fig. 2D.
Transitioning to VAET - To verify that the additional transfer pathways—provided by the slow vibrational mode—are responsible for the smooth transfer profile shown in Fig. 2D for the non-degenerate CT case, we decrease the electronic coupling strength and the vibronic coupling strength to the slow vibrational mode in Figs. 3A-C (see Appendix G for more details on VAET-CT crossover). Weaker electronic coupling ($|V|<\lambda_{1}/4$) places the system closer to the nonadiabatic regime with respect to the fast mode—a regime in which the donor and acceptor populations are localized on their respective potential energy surfaces along $y_{1}$, and resonant excitation transfers between well-defined donor and acceptor vibronic states occur at $\Delta E\approx\ell\omega_{1}$, with $\ell$ being an integer [28, 20] (see Appendix C.1). In the single-mode case, this results in the manifestation of the vibrational mode structure in the transfer rate spectrum (solid blue curve in Fig. 3A). At the same time, lesser vibronic coupling to the slow mode reduces the distortion of the potential energy surfaces along $y_{2}$, making the transfer processes involving the slow mode to enter the VAET regime and introducing additional transfer rate resonances energetically enabled by the slow mode.
As shown by the resolved resonances in Fig. 3A, the additional peaks around the nonadiabatic transfer resonances of the fast mode, which coincide with the single-mode spectrum, are induced by the second mode and located at $\Delta E\approx\ell\omega_{1}+\omega_{2}$, where the $\omega_{2}$ contribution comes from the first-order VAET process bridging the energy gap between vibronic states defined by the fast mode (see the discussion of the VAET regime below). With either a larger electronic coupling strength $V$ or a stronger vibronic coupling strength to the slow mode $g_{2}$, the additional resonances caused by the presence of the slow mode merge with the fast-mode transfer resonances into a smooth spectrum, as shown, for example, in Fig. 2D. This highlights the crucial role of the simultaneous presence of fast and slow vibrational modes in reducing the CT dependence on the donor-acceptor energy gap through the additional transfer pathways.

### v3

@@ Abstract
The high degree of control and tunability of programmable quantum platforms—such as trapped ions [8], superconducting qubits [12], and photonic devices [13]—potentially provides an alternative approach to large-scale classical computations for studying condensed-phase chemical dynamics through direct quantum simulation. In particular, trapped-ion analog and analog-digital simulators have already realized a variety of chemical dynamics with fully programmable system parameters and accurate time-resolved features [14, 15, 16, 17, 18]. With engineered reservoirs on their native spin and bosonic degrees of freedom, trapped-ion systems have also been recently used to study non-equilibrium quantum reaction dynamics [19, 20, 21, 22, 23].
@@
In this work, we present the first trapped-ion simulation in which unitary spin–phonon couplings and mode-selective dissipation via cooling are independently programmed, enabling real-time open-system emulation of excitation transfer dynamics in a two-mode LVCM. Engineering spin-phonon systems with multiple bosonic modes, combined with environment engineering, represents an important step toward realizing models of molecular systems in which multiple intramolecular (fast) modes coexist with long-wavelength, low-frequency environment-mediated modes—both of which play crucial roles in excitation-energy transfer [24]. Here, by using both ground-state and optical qubits, we achieve simultaneous control over mode frequencies, vibronic couplings, and dissipation rates. We characterize the low-temperature transfer rate across a wide parameter range, highlighting the roles of mode degeneracy and vibronic coupling strength. We focus on two transfer regimes (see Fig. 1A): charge transfer (CT) and vibrationally assisted exciton transfer (VAET), which correspond to strong and weak vibronic coupling, respectively. The distinction is made based on the differing phenomenological effects of the vibronic coupling on typical excitation and charge transfer processes [6, 25]. When the vibrational modes are strongly coupled to the electronic states of the system, they reshape the potential energy landscape and actively influence the charge transfer reaction [26]. By contrast, weak couplings between the vibrational and electronic degrees of freedom primarily allow phonons to assist coherence between the electronic eigenstates [27]. We note that charge transfer can also occur in the weak vibronic coupling regime [28]. However, in this study, we use this term to refer specifically to the strong vibronic coupling regime, with which it is more commonly associated.
Figure 1: Two-mode LVCM with engineered reservoirs. (A) Schematic diagram illustrating two regimes of transfer dynamics, defined by vibronic coupling strengths. In the VAET regime, the harmonic oscillators are weakly coupled to the donor–acceptor electronic system, whereas the strong vibronic couplings in the CT regime introduce significant displacements, resulting in distinct two-dimensional potential energy surfaces for the donor (red) and acceptor (blue) states. The electronic coupling opens a $2V$ avoided crossing between these surfaces. For clarity, we show only the $y_{2}=0$ cut across the two-dimensional potential energy surfaces, with the vibrational levels associated with the fast $(\omega_{1})$ and slow ($\omega_{2}$) modes, represented by their wavefunctions and horizontal lines, respectively. In both regimes, the vibrational modes undergo dissipation, represented by the wiggly arrows. (B) Top (dashed panel): Illustration of a first-order VAET process at zero temperature, where weak vibronic coupling coherently drives transitions between the eigenstates of the electronically coupled system via a single-excitation exchange with a vibrational mode (see Appendix C.2). Specifically, when energy is released from the higher-energy eigenstate, the vibrational mode gains a phonon excitation, which subsequently dissipates into the environment. Bottom (dashed panel): Two constructively interfering transfer pathways underlying the second-order VAET process, which bridges the energy gap of $\sqrt{\Delta E^{2}+(2V)^{2}}=\omega_{1}+\omega_{2}$. (C) Top: Two-dimensional energy landscape in the CT regime, with coordinates defined by the two vibrational modes. The blue and pink surfaces correspond to the system’s upper and lower adiabatic surfaces under strong electronic coupling, respectively. Bottom: Contour plot of the lower adiabatic surface, with the green arrow indicating the direction of charge transfer along the effective reaction coordinate.
@@
[…]
@@
In this model, the charge transfer (CT) regime is realized when the vibronic coupling is comparable to or larger than the harmonic frequency ($g_{i}\gtrsim\omega_{i}$, also known as strong coupling in quantum optics), which is characteristic of many situations involving charge motion in chemical and biological reactions, ranging from redox catalysis to solvent-induced electronic delocalization [30, 31, 32, 26]. The donor and acceptor electronic sites in this regime are described by uncoupled two-dimensional potential energy surfaces with respect to the ($y_{1}$, $y_{2}$) spatial coordinates, where $y_{i}=y_{i0}(a_{i}+a_{i}^{\dagger})/2$ with $y_{i0}=\sqrt{1/2m\omega_{i}}$ and $m$ being the particle mass. The electronic coupling $V$ mixes the two potential energy surfaces by opening an avoided crossing with a gap of $2V$ at their intersection. A strong vibronic coupling ($g_{i}\gtrsim\omega_{i}$) distorts the energy landscape by inducing a displacement of $g_{i}/\omega_{i}$ between the donor and acceptor potential energy surfaces along $y_{i}$, providing the electronic coupling with a dependence on the overlaps of the displaced oscillator states [33, 26, 20]. During CT, the donor population undergoes crossing of the energy barrier along an effective reaction coordinate, roughly defined by the axis $y_{1}+y_{2}$ in the degenerate ($\omega_{1}=\omega_{2}$) case. The total reorganization energy of the system is defined as $\lambda=\sum_{i}\lambda_{i}\equiv\sum_{i}g_{i}^{2}/\omega_{i}$, which is the energy required to displace the wave packet on an uncoupled potential energy surface by the distances of $|g_{1}/\omega_{1}|$ and $|g_{2}/\omega_{2}|$ along the $y_{1}$ and $y_{2}$ axes, respectively. When the electronic coupling is small ($V\approx 0$), the reorganization energy $\lambda$ and energy difference between the two surfaces $\Delta E$ determine the classical activation energy, defined by $U=(\Delta E+\lambda)^{2}/4\lambda$ [26, 20].
Within the CT regime, there exist two electronic coupling regimes: the strictly nonadiabatic (or perturbative) regime with $|V|<\lambda_{i}/4$ and the strongly adiabatic (or non-perturbative) regime with $|V|\sim\lambda_{i}/4$ [26, 20, 34]. In the former, Fermi’s golden rule—treating the electronic coupling as a perturbation to the uncoupled donor-acceptor system—predicts that transfer dynamics occur when the quantized vibronic energy levels of the donor and acceptor surfaces match, resulting in transfer rate resonances at:
$\Delta E\approx\ell_{1}\omega_{1}+\ell_{2}\omega_{2},$ (2)
with $\ell_{1}$ and $\ell_{2}$ being integers (see Appendix C.1). On the other hand, at larger electronic coupling ($|V|\sim\lambda_{i}/4$), the avoided crossing becomes appreciable, leading to the hybridization of the two-dimensional donor and acceptor potential energy surfaces into upper and lower adiabatic energy surfaces (see Fig. 1C). This results in delocalized eigenstates that are superpositions of donor and acceptor vibronic states [26, 20]. In this regime, Fermi’s golden rule no longer applies, which provides strong motivation to study these transfer conditions experimentally.
Figure 2: Trapped-ion quantum simulation of a two-mode LVCM with engineered reservoirs. (A) Experimental setup for studying open-system LVCM dynamics with two vibrational modes using a 171Yb+ -172Yb+ ion chain, whose motional degrees of freedom are shared (represented by the connecting spring). Insets: simplified level schemes for 171Yb+ and 172Yb+ qubits. Stimulated Raman transitions on the 171Yb+ ground-state qubit with 355 nm beams (purple) are used to engineer the Hamiltonian in Eq. (1). The optical qubit of 172Yb+ is addressed with a 435.5 nm beam (light blue) and a 935 nm repumper beam (red line in the 172Yb+ inset) for sympathetic cooling. (B) Experimental sequence used to measure the time-resolved evolution of the excitation at the donor electronic site (see Methods for details). Each dashed red line indicates the state of the system after each preparation pulse.
@@
Conversely, weak vibronic coupling ($g_{i}\ll\omega_{i}$) is more characteristic of vibrationally assisted exciton transfer (VAET), as occurs between pigments in light-harvesting compounds and their reaction centers [14]. In this regime, excitation transfer between donor and acceptor sites with an energy separation $\Delta E$ is enabled by the electronic coupling term $V\sigma_{x}$, resulting in two eigenenergies of the total electronic system that differ by $\sqrt{\Delta E^{2}+(2V)^{2}}$. Unlike in the CT regime, where strong vibronic couplings displace the donor and acceptor energy surfaces and thereby define the vibronic states that participate in excitation transfer, the couplings between the electronic system and the vibrational modes in the VAET regime are weak. From Fermi’s golden rule, these weak couplings perturbatively facilitate quantized energy exchange between the oscillators and the electronically coupled excitation sites, leading to transfer resonances at:
$\Delta E\approx\sqrt{(\ell_{1}\omega_{1}+\ell_{2}\omega_{2})^{2}-(2V)^{2}},$ (3)
with $\ell_{1}$ and $\ell_{2}$ being integers. At resonance, the combined vibrational energy $\ell_{1}\omega_{1}+\ell_{2}\omega_{2}$ provided by the two modes exactly bridges the electronic energy gap, thereby assisting the transfer [14, 21, 27] (see Appendix C.2).
Figure 3: Transfer rates in the CT regime. (A) Transfer rate spectrum of degenerate CT ($\omega_{1}\!=\!\omega_{2}\!\equiv\!\omega$) with $(V,g_{1},\gamma_{1},g_{2},\gamma_{2})=(0.200,1.200,0.036,1.100,0.040)\omega$. Red circles represent experimental data with error bars estimated via bootstrapping (see Methods). The solid red curve shows the transfer rate calculated from Eq. (4), using the definition in Eq. (5) and including spin decoherence ($\gamma_{z}\!=\!0.0014\omega$) and motional dephasing of both modes ($\gamma_{im}\!=\!0.0160\omega$, with $i=1,2$). The blue curve shows the numerical result for the single-mode CT case, where $\omega_{2}\!=\!g_{2}\!=\!\gamma_{2}\!=\!\gamma_{2m}\!=\!0$. (B and C) Experimental and numerical donor population evolution $P_{D}(t)$ versus energy gap $\Delta E$ and the number of vibrational oscillations $\omega t/2\pi$, using the same parameters as in the red circles and solid red curve in (A), respectively. Here, the detunings from the two tilt modes for encoding the degenerate vibrational energies are both set to $\delta_{1}=\delta_{2}=-2\pi\times 5$ kHz. (D to F) Same layout as in (A to C) for the non-degenerate CT case ($\omega_{1}\!>\!\omega_{2}$) with $(V,g_{1},\gamma_{1},\omega_{2},\gamma_{z},\gamma_{im})=(0.138,1.029,0.023,0.375,0.0009,0.0100)\omega_{1}$ and $(g_{2},\gamma_{2})\!=\!(0.993,0.027)\omega_{2}$. The characteristic timescale is defined by the fast vibrational frequency $\omega_{1}$. In this case, the detuning from the inner (outer) tilt mode, which encodes the fast (slow) vibrational energy, is set to $\delta_{1}\!=\!-2\pi\!\times\!8$ kHz ($\delta_{2}\!=\!-2\pi\!\times\!3$ kHz).
@@
In most realistic chemical situations, the interaction between the donor-acceptor vibronic system and the external environment causes the vibrational modes to undergo incoherent dissipation. Thus, the full Hamiltonian, $H_{\rm total}=H+H_{\rm b}+H_{\rm sb}$, also includes the bath degrees of freedom $H_{\rm b}$, described by a large collection of continuous harmonic oscillators, as well as a linear coupling between the bath and the system’s vibrational modes, given by $H_{\rm sb}$. The correlations of the environment and their influence on the system can be characterized by a continuous spectral density function $J(\omega)$. Under the assumptions of a Markovian environment with Ohmic spectral densities [35, 26] and vibrational dissipation rates that are weaker than the vibrational and thermal energies ($\gamma_{i}\ll\omega_{i},\gamma_{i}\ll k_{B}T_{i}$) [36], the dissipative dynamics of the system can be effectively described by a Lindblad master equation:
@@
Our experiment is performed in a driven rotating frame, where we use two $\pi/2$ pulses to map the $z$ spin basis in Eq. (1) onto the $y$ basis. This method allows us to use the ion-light interactions between the 171Yb+ qubit and the 355 nm Raman laser tones to independently engineer individual terms of the Hamiltonian in Eq. (1) with precise control. Two laser tones resonant with the qubit frequency generate the single-qubit operations ($V\sigma_{x}$ and $({\Delta E}/{2})\,\sigma_{z}$ terms), while two other laser tone pairs at frequencies $\pm\mu_{i}=\pm(\omega_{\text{tilt},i}+\delta_{i})$ from the qubit resonance realize the vibronic coupling and harmonic terms in Eq. (1). Here, $\delta_{i}$ is the detuning from its respective radial tilt mode at frequency $\omega_{\text{tilt},i}$, which determines the vibrational energy $\omega_{i}$ through $\delta_{i}\equiv-\omega_{i}$ [37, 20].
[…]
In this work, we focus on the non-perturbative quantum regime [26], where the electronic coupling strength is strong ($|V|\!\sim\!\lambda_{i}/4$), and the bath temperatures are low ($\bar{n}_{i}\sim 0.1$–0.2, see Methods). We compare the excitation transfer behaviors of the two-mode systems with those of their single-mode counterparts, which have been experimentally realized on trapped-ion simulators in Refs. [20, 14, 21] to various extents. In the following, we characterize the transfer dynamics by measuring the inverse lifetime of the donor population, $P_{D}=(\langle\sigma_{z}\rangle+1)/2$, which is described by [33, 26, 20, 34]:
$k_{T}=\frac{\int P_{D}(t)dt}{\int tP_{D}(t)dt}.$ (5)
@@
Charge transfer (CT) - We first study non-perturbative CT with $g_{i}\sim\omega_{i}$ in two different cases: degenerate ($\omega_{1}=\omega_{2}\equiv\omega$) and non-degenerate ($\omega_{1}>\omega_{2}$), where the approximate analysis breaks down. In the degenerate case, we observe in Fig. 3A that, in the two-mode model, the exothermic region characterized by monotonically increasing transfer rates with respect to the energy offset $\Delta E$, caused by both the broadening effect of strong electronic coupling to off-resonant states and by the dissipation rates that limit the transfer rates ($|V|>\gamma_{i}$) [26, 20], extends to a higher energy gap, $\Delta E\approx 4\omega$, compared to its single-mode counterpart ($\Delta E\approx 3\omega$). This extension can be attributed to the larger number of state configurations available when two vibrational modes (two dimensions) are present, rather than just one (one dimension). We note that it is not possible to adjust the Hamiltonian parameters of a single-mode model to reproduce this observation.
@@
Moreover, we remark that resonant peaks at $\Delta E\approx\ell\omega$, with $\ell$ being an integer, appear when the energy difference is sufficiently high ($\Delta E\gtrsim 5\omega$). Under this energy gap condition, the initially localized donor state has significant overlaps with the eigenstates of the upper hybridized surface, leading to population trapping, as explained in Refs. [26, 20]. At these resonant peaks, the trapped population is released from the upper hybridized surface to the lower hybridized surface during the evolution. Owing to the large energy offset and the increased number of state configurations in the two-mode degenerate case (relative to the single-mode case), the steady states of these resonant transfers exhibit greater overlaps with the acceptor states, as shown in Figs. 3B and 3C, where the final donor populations, $P_{D}(t_{\rm sim})$, are closer to zero than the steady state populations of the single-mode counterpart (see Fig. S7B in Appendix H). This results in enhanced transfer rates at large energy gaps $\Delta E\gtrsim 5\omega$, compared to the single-mode case.
@@
Conversely, when $\omega_{1}>\omega_{2}$, the donor-acceptor energy landscape supports highly delocalized states along $y_{2}$, as the system lies deeply in the adiabatic regime ($|V|>\lambda_{2}/4$) [26, 20]. This introduces multiple states delocalized along the $y_{2}$ direction, in addition to the existing delocalization along $y_{1}$ provided by the fast vibrational mode levels. Together, these enable additional transfer channels among the highly delocalized two-dimensional vibronic states across the donor-acceptor energy gap. Thus, the transfer process in the non-degenerate case is less sensitive to the donor-acceptor energy offset, making it more robust to variations in this parameter. Due to our experimental resolution (see Appendix B), we cannot resolve the resonances associated with these transfer channels and instead observe a smooth transfer rate curve in Fig. 3D. In addition, unlike the degenerate case, the transfer rates at the resonances with large $\Delta E$ are not evidently enhanced relative to those of the single-mode system. We emphasize that the non-degenerate two-mode spectrum observed in Fig. 3D cannot be obtained by tuning the parameters of a single-mode model.